I need your help with this question:
Assuming that no questions are left unanswered, in how many ways can a five-question quiz with 5-choice multiple choice questions be answered?
The quiz can be answered in ___ ways.

Answer by jim_thompson5910(35256)

For each slot, you have 5 choices. Since there are 5 slots (and you can repeat choices), this means that you have 5*5*5*5*5 = 5^5 = 3125 different ways to answer this quiz.

Answer: The quiz can be answered in _3125_ ways.